OpenStudy (anonymous):

Can someone please please help me with finding all roots Please x^4 +18x^2 +81=0 x^3 +2x^2 −15x=0 x^4 +6x^3 +8x^2 =0

OpenStudy (anonymous):

for equation no1 since the second term has positive sign, so the root will be the root of the first term plus the root of the third term

OpenStudy (aum):

x^4 + 18x^2 + 81 = 0 x^4 + 9x^2 + 9x^2 + 81 = 0 x^2(X^2 + 9) + 9(x^2 + 9) = 0 (x^2+9)(x^2+9) = 0 (x^2+9)^2 = 0 x^2 + 9 = 0 x^2 = -9 x = -3i, 3i

OpenStudy (anonymous):

So would it factor to (x^2+9)2=0 then all the roots would be Roots: {3i multiplicity 2, −3i multiplicity 2}

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

Thank you so much you are the only person who continued to help me

OpenStudy (aum):

You are welcome. For the second problem, factor out x. You will be left with a quadratic which you can then factor. For the third problem, factor out x^2. You will be left with a quadratic which you can then factor.

OpenStudy (anonymous):

Thank you :)

OpenStudy (aum):

yw.

OpenStudy (anonymous):

I know this is a lot to ask and I hope I'm not bothering you but is this the answer to the 2nd one Factors to: x(x − 3)(x + 5)= 0 Roots: {0, 3, −5}

OpenStudy (aum):

Yes!

OpenStudy (anonymous):

AHHH Yes thank you so much!

OpenStudy (aum):

yw.